OpenStudy (anonymous):

A rectangular sign had a perimeter of 44 feet and a length of 14 feet. A designer shortened the length so that the sign became a square with the same width as the old sign. What was the area of the new sign? A. 64 ft2 B. 196 ft2 C. 112 ft2 D. 256 ft2

OpenStudy (anonymous):

perimeter function of length and width so p(l, w) given: p(l, w) = 44 and l = 14 need a function for p in terms of l and w p(l, w) = 2l + 2w l = 14 so plug this in p(14, w) = 2(14) + 2w p(14, w) = 28 + 2w also we know p(14, w) = 44 so 44 = 28 + 2w 44 - 28 = 2w 16 = 2w 8 = w now we want a square with width 8, so square 8 8^2 = 64 area = 64